Solve the matrix (I-A)X=D for the given matrices A,D,and X.
[0.1 0.1]=A
[0.2 0.1]
[7 ]=D
[14]
[x]=X
[y]
Please enter each number as a decimal
Round each answer to two decimals
x=
y=

Answer by stanbon(48568)

Solve the matrix (I-A)X=D for the given matrices A,D,and X.
[0.1 0.1]=A
[0.2 0.1]
------------------
[7 ]=D
[14]
-------------
[x]=X
[y]
---------------------
Please enter each number as a decimal
Round each answer to two decimals
x= 0.1*7+0.1*14 = 2.4
y= 0.2*7+0.1*14 = 2.8
==========================
Cheers,
Stan H.